A juice drink manufacture has found mixing pineapple juice concentrate and fresh squeezed orange juice is the most cost effective way to produced its breakfast beverage. The pineapple juice comes in two concentrated forms,21% and 70% pineapple juice, which must be combined into a solution that can be mixed with the orange juice. If 5o gallons of 21% juice is used, how many gallons of the 70% juice must be used to obtain a 60% pineapple juice solution?
Answer by Paul(988)

21(5)+70x=60(x+5)
105+70x=60x+300
10x=195
x=19.5
Hence about 19.5 gallons of 70% jucie must be to obtain 60% pineaple juice.
paul.